What is the distance between points A and B if A has coordinates (-5,7) and B (1,-5)? Round to two decimal places
Answer by rfer(16322)

a^2+b^2=c^2
6^2+12^2=c^2
36+144=c^2
180=c^2
sqrt 180=13.42 units